OpenStudy (waheguru):

I know how to find finite differences such as first or second differences. But what if the x co-ordinates given are in random order.

OpenStudy (mertsj):

I don't even know what it means...much less how to solve it.

OpenStudy (waheguru):

For example x y 1 3 3 -3 4 -9 2 1

OpenStudy (waheguru):

When I find the first differences for the first y consecutive y values or even the second consecutive y values, I will not know if its first or secind differences because the x vaules are random and are not increasing by the same amount, what do I do in this case

OpenStudy (waheguru):

does tis make sense @Mertsj

OpenStudy (anonymous):

Newton's method of finite differences won't work if the \(x\)'s aren't equally spaced apart.

OpenStudy (waheguru):

There is a way but I do not recall...

OpenStudy (tkhunny):

1) Sort them. Why is that hard? 2) Points not equally spaced simply requires a different denominator. 0 1 1 5 3 9 Unit Differences are (5-1)/(1-0) and (9-5)/(3-1)

OpenStudy (waheguru):

Yes but this was a simple example, what if they were not sortable?

OpenStudy (waheguru):

I want to how how to find wether they are first or second differences without sorting them because realistically, we cannot always sort them

OpenStudy (anonymous):

ill say it again, As I have said this @tkhunny guy/ girl is good

OpenStudy (tkhunny):

Ummm... Real Numbers are Well-Ordered. Why are the not sortable? Realistically, we can always sort them.

OpenStudy (waheguru):

Ok, say the pattern was 1 2 3 6 4 12 7 18 How can we sort this, and find if they are first or second differences

OpenStudy (tkhunny):

I suppose we could have precision problems so that we can't actually tell if they are sorted. I suppose, if they cannot be sorted, that we can still index them. This would be essentially "sorted". Oh, how I wish I understood the question. What are these? 1 2 3 6 4 12 7 18 Are you saying these are NOT abscissas and ordinates? We believe these to be differences of some kind?

OpenStudy (waheguru):

These are co-ordinates the first are x and second are y

OpenStudy (tkhunny):

Okay, how are they not sorted?

OpenStudy (waheguru):

we are missing x values of 2 5 and 6..

OpenStudy (tkhunny):

Well, that seems to have nothing to do with sorting. They are just not evenly-spaced. This is where I started this conversation. You must unitize the differences. 1 2 3 6 4 12 7 18 Available 1st Differences (6-2)/(3-1) = 4/2 = 2 (12-6)/(4-3) = 6/1 = 6 (7-4)/(18-12) = 3/6 = 1/2 Available 2nd Differences (6-2)/(4-1) = 4/3 ((1/2)-6)/(7-4) = (-11/2)/3 = -11/6 These are not quite the Unit Differences you are used to. They are more of an average over the provided range.

OpenStudy (waheguru):

Wow, but would that would for x y 1 3 3 -3 4 -9 2 1

OpenStudy (tkhunny):

You can see these date from y = x^3, one with complete data and the other with some values missing. 1 1 2 8 7 3 27 19 12 4 64 37 18 6 5 125 61 24 6 6 216 91 30 6 7 343 127 36 6 1 1 3 27 13 4 64 37 8 7 343 93 14 1 It's not as good as one might want, but it's more than nothing.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Divided_differences